### v6

# Worst-case search in constrained uncertainty space for robust $\mathcal{H}_{\infty}$ synthesis

Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of standard, readily available smooth optimization to address this nonsmooth constrained optimization problem. Specifically, we prove that for such functions, any subgradient satisfy Karush-Kuhn-Tucker (KKT) conditions at a local minimum, and that any accumulation point of the sequential quadratic programming (SQP) is a KKT point. From a practical point of view, we combine this local exploitation with a global exploration using Monte-Carlo sampling. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. The proposed approach is illustrated through the robust control of a mechanical system. We show that this method enables fast detection of rare worst-case configurations, and that the robust controller optimization converges with a limited number of active configurations.
keywords:
Robust control, uncertain systems, nonsmooth optimization
thanks: Corresponding author.
1 Introduction
The $\mathcal{H}_{\infty}$ norm is a widely used metric to quantify the performance of control systems. For a linear system, it is equivalently defined as the induced 2-norm, hence representing the maximum energy amplification from input to output, or as the maximum peak in the frequency domain:
[…]
where $T_{zw}(\mathrm{s})$ is a transfer function between input $w$ and output $z$. Another system norm is the $\mathcal{H}_{2}$ norm, representing the energy of the impulse response, or the output’s standard deviation in response to a white noise of unit power spectral density. It is thus natural to seek to optimize these norms in control problems. Moreover, practical problems include parametric uncertainties that may severely degrade the performance and the stability if not addressed during controller optimization. Consequently, the standard robust structured $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ control problem is formulated as the min-max optimization problem:
[…]
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and the transfer function $T_{zw}(\mathrm{s},K,\delta)$ is now written as a function of $K$ the controller to be optimized, $\mathcal{K}$ is the set of controllers of the chosen structure, and $\delta$ is the vector of $k$ uncertain parameters in a parametric space $\mathcal{D}$. Multiple control requirements are generally addressed with the following formulation:
[…]
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters. In this context, an overview of this control problem and of the modern resolution method, relying on nonsmooth optimization, is presented in [9]. The goal of this paper is to address the presence of nonlinear constraints restricting this hypercube.
We mainly focus on the inner maximization problem, often called worst-case search; solving it will then allow us to treat Problems (1) and (2) by leveraging state-of-art nonsmooth optimization of the controller [8]. Namely, the worst-case search in performance:
[…]
and the worst-case search in stability, implicitly required to ensure finiteness of the norms, treating it as in [3]:
$\underset{\delta\in\mathcal{D}}{\max}\;\alpha(A(K,\delta))$ (4)
where $A(K,\delta)$ is the matrix A of the state-space representation of the closed loop, and $\alpha(\cdot)$ denotes the spectral abscissa.
When the parametric space is the whole hypercube $1,1]^{k}$, Matlab’s function systune, relying on the non-smooth optimization of [8, 3], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [20] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [15] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. The $\mu$-analysis [26] was developed for this purpose, generally based on branch-and-bound algorithms [34]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [12, 36].
This paper addresses the case where nonlinear constraints, noted $c_{i}(\delta)\leq 0$ and possibly including equality constraints, restrict the hypercube:
$\mathcal{D}=\{\delta\in[-1,1]^{k}\text{ s.t. }c_{i}(\delta)\leq 0,\;1\leq i\leq n\}\;$ (5)
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Our focus is twofold. Firstly, we seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve Problems (1) and (2) in reasonable time. Secondly, we aim to assess their reliability in consistently detecting global worst-case configurations as a validation tool.
Problem (3) is widely seen as a non trivial task, due to the nonsmoothness of the $\mathcal{H}_{\infty}$ norm. Nevertheless, [3] links the $\mathcal{H}_{\infty}$ norm, as a function of a structured controller or uncertainty, to the theory of lower- and upper-$C^{1}$ functions [37], a class of functions with a variety of regularity properties [33]. Multiple references suggest that the minimization of upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In [3], the authors show the convergence of their bundle method even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization) when applied to such functions. This work follows [24] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [10] in the context of robust control. In [25], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationary points when applied to upper-$C^{1}$ functions (in the context of unconstrained optimization).
In this paper, we demonstrate new properties of the class of upper-$C^{1}$ functions, and in particular that any subgradient verifies Karush-Kuhn-Tucker (KKT) conditions when the constraint functions $c_{i}$ are differentiable. We also propose a slight adaptation of the Sequential Quadratic Programming (SQP) [28], a smooth constrained optimization algorithm, to the upper-$C^{1}$ case, and demonstrate the convergence to KKT points. We deduce that this algorithm can be applied to the nonsmooth problem (3), and to (4) under an additional assumption. From a practical point of view, we also take advantage of Monte-Carlo sampling and its stochastic properties [38] to perform a global exploration of the parameter space (in the stochastic sense), with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is combined with multi-models controller optimization [8] to perform robust control, with the possibility to add multiple worst-case configurations to the active set at each iteration. The proposed approach is illustrated through the control of a realistic mechanical system, which typically exhibits constraints of the form (5).
We emphasize that the proposed approach, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters, since the exact computation of the structured singular value is NP-hard [39], forcing reliance on bounds that are often either intractable to compute or overly conservative. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), albeit without formal guarantee of global optimality. Furthermore, the nonsmooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration in the stochastic sense; it also opens the possibility to speed up the optimization by adding multiple active configurations at each iteration.
Notations
Given $T_{zw}(s)$ a transfer function between an input $w$ and an output $z$, $T_{zw}(s)||$ may denote either the $\mathcal{H}_{\infty}$ or the $\mathcal{H}_{2}$ norm for the sake of generality; however, the paper will mostly focus on the $\mathcal{H}_{\infty}$ norm (because the $\mathcal{H}_{2}$ norm is differentiable, and thus does not cause any particular difficulty). $\alpha(\cdot)$ denotes the spectral abscissa:
$\alpha(M)\vcentcolon=\max\{\textrm{Re}(\lambda):\lambda\text{ is an eigenvalue of }M\}\;.$
Given $f:\mathbb{R}^{n}\to\mathbb{R}$ a locally Lipschitz function, the Clarke subdifferential of $f$ at point $x$ is denoted $\partial f(x)$. When defined, the directional derivative of $f$ at $x$ in direction $d$ is defined as
$f^{\prime}(x,d)=\lim_{t\to 0^{+}}\frac{f(x+td)-f(x)}{t}\;.$
2 Constrained optimization of upper-$C^{1}$ functions
2.1 Properties of upper-$C^{1}$ functions
In this section, we introduce the class of upper-$C^{1}$ functions and demonstrate favorable properties for their minimization. As we shall see in Section 3 in the context of robust $\mathcal{H}_{\infty}$ control, the worst-case search in performance (3), and in stability (4) under an additional assumption, can be addressed with such functions.
Definition 1.
: Lower- and upper-$C^{1}$ functions [37]
A locally Lipschitz function $f:\mathbb{R}^{n}\to\mathbb{R}$ is upper-$C^{1}$ at $x\in\mathbb{R}^{n}$ if there exists a compact set $S$, a neighborhood $U$ of $x$, and a function $F:U\times S\to\mathbb{R}$ such that $F$ and $\nabla_{x}F$ are jointly continuous in $x$ and $s$, and such that
$f(x)=\min_{s\in S}F(x^{\prime},s)\quad\text{for all }x^{\prime}\in U.$
We note $I(x)$ the subset of $S$ that defines the active functions $F(\cdot,s)$ at $x$, i.e.:
$I(x)=\arg\min_{s\in S}\;F(x,s)=\{s\in S,f(x)=F(x,s)\}\;.$
$f$ is lower-$C^{1}$ if f-f is upper-$C^{1}$, i.e. $f$ can be written as a maximum, similarly to the expression above.
In this whole section, we consider $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ an upper-$C^{1}$ function. We recall that the directional derivative always exists for an upper-$C^{1}$ function [33]. Let us first consider the unconstrained problem:
$\underset{x\in\mathbb{R}^{n}}{\text{minimize}}\;f(x)$ (6)
Proposition 2.
: Descent directions and optimality of upper-$C^{1}$ functions – Unconstrained case
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$.
[…]
b) Optimality. If $x$ locally minimizes $f(x)$, then $g_{x}||=0$ for any subgradient $g_{x}\in\partial f(x)$.
{pf}
a) [25, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
$f^{\prime}(x,d)=\min_{g\in\partial f(x)}g^{T}d=\min_{u\in I(x)}\nabla_{x}F(x,u)^{T}d$ (7)
[…]
Proposition 2.a) shows that when a smooth optimization algorithm is provided with any subgradient as if it were the gradient of a smooth function, progress can still be achieved at points of nonsmoothness. Naturally, the selected direction is not necessarily the steepest descent. Proposition 2.b) ensures that an optimization algorithm can rely on the same optimality conditions as in the smooth case when picking any subgradient (we recall that, for the general case of nonsmooth functions, the optimality condition is $0\in\partial f(x)$ which is a weaker condition, also called Clarke-stationarity, and requires a specific treatment in nonsmooth optimization algorithms).
We shall now demonstrate similar results for the constrained problem:
$\begin{array}[]{lll}&\underset{x\in\mathbb{R}^{n}}{\text{minimize}}&f(x)\\ &\text{subject to }&c_{i}(x)\leq 0\;,\;\;1\leq i\leq n\end{array}$ (8)
where the $c_{i}$ are differentiable, and $f$ is upper-$C^{1}$. We do not consider equality constraints, as they can be written equivalently as two inequality constraints. The Karush-Kuhn-Tucker (KKT) conditions of Problem (8) read: there exists $g_{x}\in\partial f(x)$ and $(u_{i})_{1\leq i\leq n}$ such that:
$\begin{array}[]{l}g_{x}+\sum_{i=1}^{n}u_{i}\nabla c_{i}(x)=0\\ \forall i,\;u_{i}\geq 0,\;u_{i}c_{i}(x)=0,\;c_{i}(x)\leq 0\end{array}$ (9)
We also define the quadratic problem $Q(x,H,g_{x})$:
$\begin{array}[]{lll}&\underset{p\in\mathbb{R}^{n}}{\text{minimize}}&g_{x}^{T}p+\frac{1}{2}p^{T}Hp\\ &\text{subject to }&c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0\;,\;\;1\leq i\leq n\end{array}$ (10)
where $H$ is a symmetric definite positive matrix and $g_{x}\in\partial f(x)$. The quadratic problem $Q(x,H,g_{x})$ can be seen as a local approximation of the Problem (8) where the matrix $H$ may contain information about the curvature of both objective and constraint functions, usually serving as an approximation of the Lagrangian. $Q(x,H,g_{x})$ is originally introduced in methods such as [27] where $f$ is differentiable and $\nabla f(x)$ replaces $g_{x}$.
Finally, for $r\geq 0$, we introduce the merit function $\theta_{r}$ as in [27]:
$\theta_{r}(x)=f(x)+r\sum_{i=1}^{n}c_{i}(x)_{+}$ (11)
where $c_{i}(x)_{+}=\max(0,c_{i}(x))$.
Proposition 3.
Descent directions and optimality of upper-$C^{1}$ functions – Constrained case
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$. Let the $c_{i}$ be continuously differentiable.
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$ with $g_{x}\in\partial f(x)$ any subgradient. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$.
b) Optimality. If $x$ locally minimizes $f(x)$ subject to $n$ constraints $c_{i}(x)\leq 0$ (Problem (8)), then any subgradient $g_{x}\in\partial f(x)$ verifies the following KKT conditions: there exists $(\mu_{i})_{1\leq i\leq n}$ such that:
$\begin{array}[]{l}g_{x}+\sum_{i=1}^{n}\mu_{i}\nabla c_{i}(x)=0\\ \forall i,\;\mu_{i}\geq 0,\;\mu_{i}c_{i}(x)=0,\;c_{i}(x)\leq 0\end{array}$
{pf}
a) Cf. Appendix. The proof is a mere adaptation of [27, Theorem 3.1] to the upper-$C^{1}$ case.
b) Let us note $\mathcal{C}=\{x,c_{i}(x)\leq 0\;\forall i\}$. Since $x$ is a local minimizer of $f(x)$ in $\mathcal{C}$, there exists a ball $B(x,r_{1})$ of radius $r_{1}>0$ and of center $x$ such that $f(x^{\prime})\geq f(x)$ for all $x^{\prime}\in B(x,r_{1})\cap\mathcal{C}$. Recall that $U$ is the neighborhood of $x$ of Definition 1; as $B(x,r_{1})$ and $U$ are both non-empty neighborhoods of $x$, we can define $r_{2}>0$ such that $B(x,r_{2})\subset B(x,r_{1})\cap U$. Let us consider any of the functions $F(\cdot,s)$. For all $x^{\prime}\in B(x,r_{2})\cap\mathcal{C}$, we have $F(x^{\prime},s)\geq f(x^{\prime})\geq f(x)$, where the first inequality follows from $x^{\prime}\in U$, and the second from $x^{\prime}\in B(x,r_{1})\cap\mathcal{C}$. Assuming additionally that $F(\cdot,s)$ is active, i.e. $s\in I(x)$, then $F(x,s)=f(x)$ and we conclude that $F(x^{\prime},s)\geq F(x,s)$ for all $x^{\prime}\in B(x,r_{2})\cap\mathcal{C}$.
This shows that $x$ locally minimizes all active functions $F(\cdot,s)$ under the constraints $c_{i}(x)\leq 0$. Therefore, these functions verify the following KKT conditions: for all $s\in I(x)$, there exists $(\lambda_{i}(s))_{1\leq i\leq n}$ such that:
$\begin{array}[]{l}\nabla_{x}F(x,s)+\sum_{i=1}^{n}\lambda_{i}(s)\nabla c_{i}(x)=0\\ \forall i,\;\lambda_{i}(s)\geq 0,\;\lambda_{i}(s)c_{i}(x)=0,\;c_{i}(x)\leq 0\end{array}$
As shown in [20, 25], the subdifferential of an upper-$C^{1}$ function $f$ at $x$ is the convex hull formed by the gradients of the active functions $F(\cdot,s)$:
$\partial f(x)=\operatorname{co}\left\{\nabla_{x}F(x,s):s\in I(x)\right\}$
Thus, $g_{x}\in\partial f(x)$ can be written as:
$g_{x}=\sum_{j=1}^{p}a_{s_{j}}\nabla_{x}F(x,s_{j})\;,\;\;s_{j}\in I(x)\;\forall j$
where the $a_{s_{j}}\geq 0$ are such that $\sum_{j=1}^{p}a_{s_{j}}=1$ (and $p$ is at most the number of parameters plus one). Then, we have:
$g_{x}$ $=-\sum_{j=1}^{p}a_{s_{j}}\sum_{i=1}^{n}\lambda_{i}(s_{j})\nabla c_{i}(x)$
$=-\sum_{i=1}^{n}\left(\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})\right)\nabla c_{i}(x)$
$=-\sum_{i=1}^{n}\mu_{i}\nabla c_{i}(x)$
where $\mu_{i}:=\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})$. For all $i$: $\mu_{i}\geq 0$ (because all $a_{s_{j}}$ and $\lambda_{i}(s_{j})$ are $\geq 0$), and $\mu_{i}c_{i}(x)=0$ (because if $c_{i}(x)\neq 0$, then all $\lambda_{i}(s_{j})$ are 0). This proves Proposition 3.b).
Remark 4.
Proposition 2.b) is also proved in [25, Proposition 4]. Proposition 3.a) is a generalization of [27, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. It will be used in next section to prove convergence of the optimization algorithm. To the best of our knowledge, Proposition 3.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
2.2 Sequential quadratic programming (SQP) algorithm for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm adapted for upper-$C^{1}$ functions to solve Problem (8). It is described in Algorithm 1, and is identical to that of the original reference [27] for differentiable functions, except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (Eq. (10)). We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied).
Algorithm 1 Sequential Quadratic Programming (SQP) algorithm [27] adapted for upper-$C^{1}$ functions
Initialization: $x_{0}\in\mathbb{R}^{n}$, $H_{0}\in\mathbb{R}^{n\times n}$ positive definite, $r>0$, $\lambda_{max}>0$.
Step 1: Pick any $g_{x_{k}}\in\partial f(x_{k})$. Find a KKT point $p_{k}$ of the quadratic problem $Q(x_{k},H_{k},g_{x_{k}})$.
Step 2: Set:
$x_{k+1}=x_{k}+\lambda_{k}p_{k}$
and find a $\lambda_{k}\in[0,\lambda_{max}]$ satisfying:
$\theta_{r}(x_{k+1})\leq\min_{0\leq\lambda\leq\lambda_{max}}\theta_{r}(x_{k}+\lambda p_{k})+\epsilon_{k}$
where $\theta_{r}$ is defined in Eq. (11), and ($\epsilon_{k}$) is a sequence of nonnegative numbers satisfying:
$\sum_{i=0}^{\infty}\epsilon_{i}<\infty$
Step 3: Update $H_{k+1}$ by some scheme that keeps it definite positive.
To prove the convergence of Algorithm 1, the following lemmas are necessary.
Lemma 5.
: Supermonotonicity of the subdifferential of upper-$C^{1}$ functions
Let $f$ be upper-$C^{1}$ at $x$. For every $\epsilon>0$, there exists $\delta>0$ such that for all $x_{0},x^{\prime}_{0}\in\mathcal{B}(x,\delta)$:
[…]
where $g_{x_{0}}\in\partial f(x_{0})$ and $g_{x^{\prime}_{0}}\in\partial f(x^{\prime}_{0})$.
{pf}
This comes from [37, Theorem 3.9] (submonotonicity of the subdifferential of lower-$C^{1}$ functions) and is also noted for example in [5, Lemma 6].
Lemma 6.
: Perturbation of a quadratic problem
Let us consider the two quadratic problems:
$x_{0}=\arg\min\{\;Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\;\;\\ \text{subject to }\;Gx\leq g\}$
$x_{0}^{\prime}=\arg\min\{\;Q^{\prime}(x)=\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\;\;\\ \text{subject to }\;G^{\prime}x\leq g^{\prime}\}$
[…]
$x_{0}^{\prime}-x_{0}||\leq c\epsilon$
whenever $\epsilon<\epsilon_{0}$.
{pf}
Cf. Appendix. The proof is a mere adaptation of [22].
Proposition 7 extends the convergence result of SQP established in [27, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions, when providing any subgradient where a gradient would be expected. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [27, Theorem 3.2] are considered.
Proposition 7.
: Convergence of SQP for upper-$C^{1}$ functions
Let $f$ be upper-$C^{1}$ and the $c_{i}$ be continuously differentiable. Assume that the following conditions are satisfied:
(i) there exists $\alpha,\beta>0$ such that, for each $k$ and for any $x\in\mathbb{R}^{n}$:
$\alpha x^{T}x\leq x^{T}H_{k}x\leq\beta x^{T}x$
(ii) For each $k$, there exists a KKT point of $Q(x_{k},H_{k},g_{x_{k}})$ (Eq. (10)) with a Lagrange multiplier vector $u_{k}$ such that $u_{k}||_{\infty}\leq r$ (vector $\infty$-norm).
Then, any sequence $(x_{k})$ generated by Algorithm 1 either terminates at a KKT point of the constrained optimization problem (8), or any accumulation point $\bar{x}$ with
$S^{0}(\bar{x})=\{p:c(\bar{x})+\nabla c(\bar{x})^{T}p<0\}\neq\emptyset$
is a KKT point of Problem (8).
{pf}
The complete proof is provided in Appendix, and adapts that of [27, Theorem 3.2] to the case where f is not differentiable but only upper-$C^{1}$. Here, let us simply summarize said adaptations. In the proof of [27]:
1) it is always possible to choose $x_{k+1}$ as described in Algorithm 1 thanks to [27, Theorem 3.1]. This argument is maintained thanks to Proposition 3.a), which extends [27, Theorem 3.1] from differentiable to upper-$C^{1}$ functions.
2) given $\bar{x}$ be an accumulation point of $(x_{k})$, the continuity of the gradient of $f$ ensures that $\nabla f(x_{k})$ converges to $\nabla f(\bar{x})$. Here, a similar property is maintained as follows. Since $f$ is locally Lipschitz, its subgradients $(g_{x_{k}})$ are uniformly bounded for $k$ sufficiently large, hence there exists a subsequence that converges to a vector $\bar{g}$. Without loss of generality (passing to the subsequence if necessary) we note
$g_{x_{k}}\rightarrow\bar{g}\;,$
and the upper semi-continuity of the Clarke subdifferential [19, Proposition 1.5] yields $\bar{g}\in\partial f(\bar{x})$.
3) given the convergence of the (sub)sequences ($x_{k}$), ($H_{k}$), ($g_{x_{k}}$) and ($\nabla c_{i}(x_{k})$), the existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ follows from the continuity of $\nabla f(x)$ and from [22]. This argument is maintained in the upper-$C^{1}$ case thanks to Lemmas 5 and 6, where $x_{k+1}$ and $x_{k}$ play the roles of $x_{0}$ and $x^{\prime}_{0}$, and $g_{x_{0}}$ and $g_{x^{\prime}_{0}}$ play the roles of $k$ and $k^{\prime}$.
Proposition 3.b) establishes that local minima are necessarily KKT points for any subgradient, and Proposition 7 proves the convergence to KKT points when using any subgradient. Consequently, the standard SQP algorithm can be applied to upper-$C^{1}$ functions, by providing any subgradient where the gradient is expected. At a local minimum, any subgradient will be zero and thus enable detection of the optimality by the optimization algorithm. Conversely, the SQP algorithm converges to a KKT point, meaning that at least one subgradient is zero; note that convergence to points where all subgradients are zero, is not guaranteed.
Remark 8.
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth constrained optimization methods (for example [23, 24] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in Matlab or Python toolboxes) can be applied successfully.
3 Worst-case search in constrained parametric space
3.1 Problem formulation
In this section, the controller $K$ is fixed and, thus, voluntarily omitted in the notations. For the performance problem (3), we define for the $\mathcal{H}_{\infty}$ norm of a generic transfer function $T_{zw}(s,\delta)$:
[…]
We do not address the $\mathcal{H}_{2}$ norm in this section, because it is differentiable wherever it is well defined (cf. [32, 6, 11]) and therefore does not introduce any particular difficulty. As in [3], worst-case stability is studied by maximizing the spectral abscissa of the closed-loop matrix $A(\delta)$. Problem (4) can also be written as the following constrained optimization problem:
$\begin{array}[]{lll}&\underset{\delta\in\mathbb{R}^{n}}{\text{minimize}}&a(\delta):=-\alpha(A(\delta))\\ &\text{subject to }&-1\leq\delta_{i}\leq 1\\ &&c_{i}(\delta)\leq 0\end{array}$ (13)
3.2 Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [40, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Fig. 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability or the $\mathcal{H}_{\infty}$ norm of the transfer from $w$ to $z$.
Figure 1: LFT representation of $T_{zw}(s,\delta)$
The state-space representation of $M(s)$ is written as:
$\begin{array}[]{l}\left[\begin{array}[]{c}\dot{x}\\ \hline\cr q\\ z\end{array}\right]=\left[\begin{array}[]{c|cc}A_{1}&B_{1}&B_{2}\\ \hline\cr C_{1}&D_{11}&D_{12}\\ C_{2}&D_{21}&D_{22}\end{array}\right]\left[\begin{array}[]{c}x\\ \hline\cr p\\ w\end{array}\right]\end{array}$
Proposition 9.
: Regularity of $h_{\infty}$ [15, 8, 3]
a) $h_{\infty}$ is locally Lipschitz and Clarke subdifferentiable (in the sense of [20]) everywhere in
$\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is internally stable}\}\;.$
b) $h_{\infty}$ is upper-$C^{1}$ everywhere in $\mathcal{D}_{s}$.
c) $h_{\infty}$ is differentiable in $\mathcal{D}_{s}$ when only one frequency is active and of multiplicity one; in this case, its gradient is given in Eq. (14).
Proposition 9.a) is proved in [3, Proposition 1] and follows the work of [15] and [8]. The general subdifferential expression is provided in these references, but we will not use it in this paper. b) is proved in [3, Proposition 2]. c) immediately follows from a), and the expression is also provided in [3, Remark 5]: using the same notations, let us define $T_{qw}$ and $T_{zp}$ as:
$\left[\begin{array}[]{c|c}*&T_{qw}(s,\delta)\\ \hline\cr T_{zp}(s,\delta)&T_{zw}(s,\delta)\end{array}\right]=\left[\begin{array}[]{c|c}0&I\\ \hline\cr I&\Delta\end{array}\right]\star M(s)$
with $\star$ denoting the redheffer star product. The transfer functions $T_{zw}$, $T_{qw}$ and $T_{zp}$ correspond to the transfer between the inputs $p,w$ and outputs $q,z$ of the matrix $M(s)$ as in Fig. 1, but after closing the loop with the block $\Delta$. If $\omega_{0}$ is the single active frequency, such that the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is 1, $h_{\infty}$ is differentiable and its derivatives read:
$\frac{\partial h_{\infty}}{\partial\delta_{i}}=-\mathrm{Tr}\left(\mathrm{Re}(T_{qw}(j\omega_{0},\delta)p_{\omega_{0}}q_{\omega_{0}}^{H}T_{zp}(j\omega_{0},\delta))^{T}\Delta_{i}\right)$ (14)
where $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$ associated with $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ [3].
From the results of Section 2 and Proposition 9.b), we deduce that we can use SQP (Algorithm 1) to solve Problem (12). In particular, let us define the direction $d_{\delta}$ such that its components are defined as in Eq. (14) (even when $h_{\infty}$ is not differentiable), i.e. the i-th component of $d_{\delta}$ reads:
$d_{\delta,i}=-\mathrm{Tr}\left(\mathrm{Re}(T_{qw}(j\omega_{0},\delta)p_{\omega_{0}}q_{\omega_{0}}^{H}T_{zp}(j\omega_{0},\delta))^{T}\Delta_{i}\right)$
where $\omega_{0}$ is an active frequency – but not necessarily the only one, and the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is not necessarily 1 – and $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the corresponding normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$. Then, $d_{\delta}$ is a subgradient of $h_{\infty}$ (this is straightforward from the expression of the subdifferential given in [3, Section IV-1]); by Propositions 3 and 7, it can legitimately be used in SQP as if it were the gradient of a differentiable function.
Remark 10.
Proposition 3.b) is not anecdotal in this application. Indeed, after optimization of the controller, it is frequent to have multiple active frequencies. Therefore, it is important for the optimality criterion to be valid at a nonsmooth point, even if the function will most often be differentiable during the optimization.
Remark 11.
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$ (e.g. finding a controller that minimizes the $\mathcal{H}_{\infty}$ norm) would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). Moreover, a descent direction cannot be found by picking any subgradient. For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [17], or Clarke subdifferential-based [8, 14]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
The case of the spectral abscissa is slightly less favorable than the $\mathcal{H}_{\infty}$ norm. We recall that simple eigenvalues (i.e. those of multiplicity one) are differentiable [31]. As in [3], we call ”active eigenvalues” all eigenvalues whose real part is equal to the spectral abscissa.
Proposition 12.
: Regularity of $a$ [3, 13]
a) If all active eigenvalues of $A(\delta)$ are semi-simple (i.e. their algebraic and geometric multiplicity are equal), then $a$ is locally Lipschitz and Clarke subdifferentiable at $\delta$.
b) If all active eigenvalues of $A(\delta)$ are simple at $\delta$, then $a$ is upper-$C^{1}$ at $\delta$.
c) If $A(\delta)$ has only one active eigenvalue and it is simple, then $a$ is differentiable at $\delta$. Noting $\Delta_{i}:=\frac{\partial\Delta}{\partial\delta_{i}}$, the derivative of $a$ with regard to $\delta_{i}$ reads:
$\frac{\partial a}{\partial\delta_{i}}=-\mathrm{Tr}(\Delta_{i}^{T}\mathrm{Re}((I-D_{11}\Delta)^{-1}C_{1}\\ v_{i}u_{i}^{H}B_{1}(I-\Delta D_{11})^{-1})^{T})$ (15)
where $v_{i}$ and $u_{i}$ are right and left eigenvectors associated with the active eigenvalue, normalized with $u_{i}^{H}v_{i}=1$.
Proposition 12.a) was established in [3] following the earlier work of [13]. These references build on [18] that studied the directional derivatives of the spectral abscissa. b) is also shown in [3], and simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [31]. For c), the expression (15) is straightforward from the expression of the Clarke subdifferential given in [3].
Remark 13.
Additionally, if there are two active eigenvalues and they are complex conjugate eigenvalues, $a$ is also differentiable and the gradient given above remains valid for either of the two eigenvalues [13].
Remark 14.
Although the assumption of semi-simplicity is required for the definition of the Clarke subdifferential, the numerical experiments reported in [3] show that their bundle approach, based on Clarke subdifferentials, performs well in practice even without explicitly verifying this assumption.
Remark 15.
As for the $\mathcal{H}_{\infty}$ norm, minimizing the spectral abscissa is a more complex problem; in the literature, this is addressed with bundle methods [30, 29], gradient bundle [16], gradient sampling [17], Clarke subdifferential-based [13].
In this paper, as for $h_{\infty}$, we rely on the upper-$C^{1}$ property of $a$ of Proposition 12.b) and on the results of Section 2, when active eigenvalues are simple. However, the case of non-simple active eigenvalues is admittedly a limitation of using a smooth optimization algorithm for the worst-case search in stability. Nevertheless, this limitation is partially circumvented by the Monte-Carlo exploration proposed thereafter.
3.3 Global exploration, local exploitation
Problems (12) and (13) are constrained optimization problems involving the nonsmooth functions $h_{\infty}$ and $a$, which are upper-$C^{1}$ under the assumptions of Propositions 9 and 12. As demonstrated in Section 2, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems. This constitutes the main theoretical contribution of this paper. From a practical perspective, however, we additionally investigate the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. The advantages of this approach will be illustrated in the application of Section 5.
Local exploitation with Sequential Quadratic Programming (SQP)
SPQ is described in Algorithm 1 and, as discussed in Section 2, may be used to minimize upper-$C^{1}$ functions. A common choice for the matrix $H$ of the quadratic problem (10) is the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which iteratively updates it to serve as an approximation of the Hessian matrix of the Lagrangian. It is worth noting that BFGS was observed to perform well in [13] even for a non-smooth objective function (the spectral abscissa).
Global exploration with Monte-Carlo sampling
Monte-Carlo sampling is applied as follows: randomly pick a sample $\delta$ in the unit hypercube; evaluate $c(\delta)$: if positive, discard $\delta$, otherwise, evaluate the function $a(\delta)$ (respectively $h_{\infty}(\delta)$ or $h_{2}(\delta)$); repeat until having $N$ function evaluations. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions. To leverage the local search enabled by SQP, we envision different ways to perform the Monte-Carlo sampling.
First, assume a probability $p$ that a random sample lies in the basin of attraction of the global optimum, i.e. that a gradient-based local optimization starting from this sample converges to the global optimum. If $N$ samples are drawn and all serve as initial point for an SQP run, there is a probability $1-(1-p)^{N}$ of finding the global optimum. Thus, given a target probability $p$ (reflecting the size of the basin of attraction of the global optimum that we wish to be able to detect), and a confidence level $P$ such that $1-(1-p)^{N}\geq P$, the required number $N$ of samples is
$N\geq\frac{\mathrm{ln}(1-P)}{\mathrm{ln}(1-p)}\;.$
A second possibility, in worst-case search problems, is to pick the number $N$ of function evaluations to verify:
$N\geq\frac{\textrm{ln}(\gamma)}{\textrm{ln}(1-\epsilon)}$
where $\gamma$ and $\epsilon$ serve as confidence and accuracy levels to guarantee the probability [38]:
$\mathrm{P}(\mathrm{P}(f(\delta)<f(\delta^{*})\leq\epsilon)\geq 1-\gamma$ (16)
where $\delta^{*}$ corresponds to the lowest value of the objective function $f$ (in our case: $a(\delta)$ or $h_{\infty}(\delta)$) among the $N$ samples, i.e. that $\delta^{*}$ covers a proportion $1-\epsilon$ of the distribution with a confidence of $1-\gamma$. Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Then, a local search with SQP can be performed starting from this worst case to refine the result. Alternatively, it is also possible to draw $M$ independent groups of $N/M$ samples and to initialize a SQP with the worst case of each group; the stochastic guarantee is the same as in Eq. (16) (tied to the worst case among the total of $N$ samples) but it allows to perform multiple local searches from different points of the parametric space to increase coverage. If those SQP runs can be parallelized, this procedure comes at no extra cost.
4 Robust control in constrained parametric space
This Section leverages the worst-case search of previous section, to address robust controller optimization i.e. Problems (1) and (2) in the parametric space $\mathcal{D}$ given in (5).
4.1 Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1), follows the approach proposed in [3]. It relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations and, in this paper, essentially exploits the novel worst-case search strategy introduced in previous sections. The multi-objectives formulation is tackled in Section 4.2 with minor modifications to Algorithm 2.
The set of active configurations is initialized with the nominal configuration. Step 1 optimizes the controller for all current active configurations, based on the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section 3. If a Karush-Kuhn-Tucker (KKT) point $\delta^{*}$ of the optimization problem is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 seeks to validate the design with a more exhaustive worst-case search, meaning that the number of Monte-Carlo samples can typically be larger than in Steps 2 and  3 (where faster execution is privileged) to ensure better stochastic guarantees in accordance with Eq. (16).
Algorithm 2 Robust control optimization with constrained uncertainties (largely inspired from [3])
Initialization: Initialize the set of active configurations as $\mathcal{D}_{a}=\{0\}$.
Step 1: Multi-model synthesis.
[…]
Step 2: Worst-case search in stability.
Identify one or multiple KKT point(s) for the problem: $\delta^{*}=\arg\;\max_{\delta\in\mathcal{D}}\;\alpha(A(K^{*},\delta))$ using MC+SQP.
If $\alpha(A(K^{*},\delta))>\alpha_{max}$, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1. Otherwise continue to step 3.
Step 3: Worst-case search in performance.
[…]
Step 4: Validation
Perform a more exhaustive worst-case search in stability and performance using MC+SQP. If a configuration $\delta^{*}$ not respecting criteria of step 2 or 3 is found, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1. Otherwise, terminate.
4.2 Multi-objective formulation
Algorithm 2 is presented to address the standard problem formulation (1) involving a single $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ criterion. We note that this formulation also directly addresses the following multi-objective problem:
[…]
since it is equivalent to a single objective problem:
[…]
More generally, we consider Problem (2). The controller optimization [8] of Step 1 treats it as described in in [4]. In our worst-case performance searches (in Steps 3 and 4), we use the normalization of Eq. (17), such that the performance degradation test of Step 3 becomes, in the multi-objectives case:
$\text{max \;}\left\{\frac{1}{(1+\epsilon_{1})\beta}\underset{\delta\in\mathcal{D}}{\text{max \;}}||T_{z_{1}w_{1}}(s,K,\delta)||\;,\right.\\ \left.\frac{1}{(1+\epsilon_{2})}\;\underset{2\leq i\leq n}{\text{max \;}}\underset{\delta\in\mathcal{D}}{\text{max \;}}||T_{z_{i}w_{i}}(s,K,\delta)||\right\}>1$ (17)
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and where $\beta$ is the soft performance obtained in Step 1:
[…]
i.e. the algorithm has converged when the worst-case search degrades the soft requirement by less than $1+\epsilon_{1}$, and the hard requirements are less than $1+\epsilon_{2}$.
4.3 Adding one vs multiple KKT points to active set
Algorithm 2 gives the possibility to add one or multiple KKT points to the active configurations after a worst-case search is ran. This paragraph elaborates on how to select which KKT points to add when there are multiple available, e.g. after running several SQP in parallel.
Consider two KKT points $\delta_{1}^{*}$ and $\delta_{2}^{*}$. If their distance $\delta_{1}^{*}-\delta_{2}^{*}||_{2}$ is inferior to a predefined threshold, it is not necessary to add both of them. This threshold can typically be chosen as a fraction of the diagonal of the hypercube $2\sqrt{n}$, where $n$ is the number of uncertain parameters. However, this approach may not be very robust to the presence of uninfluential parameters, which may have very different values in $\delta_{1}^{*}$ and $\delta_{2}^{*}$ and thus contribute to the distance, without actually justifying to include both configurations to the active set.
We circumvent this limitation by weighting the distance with global sensitivity indices $w_{i}$, computed for each parameter. Such methods generally only require a few hundreds of samples, and thus come at limited extra cost. Then, the distance is defined as $W(\delta_{1}^{*}-\delta_{2}^{*})||_{2}$ with $W=\text{diag(}w_{i}$). The threshold may now be chosen as a fraction of 2w2||w|| the diagonal of the rescaled hyperrectangle. For example, in the numerical application of Section 5.4, we used the activity scores of [21], a global sensitivity metric derived from the active subspace method.
5 Illustration: control of an uncertain mechanical system
The benchmark used in this study is designed to be realistic in terms of complexity (number of states, uncertain parameters, and decision variables), challenging regarding robustness, and includes a nonlinear constraint on the uncertain parameters that is typical of mechanical systems. The tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM, on Matlab R2025a. Computation times are provided for reference, but they may vary depending on hardware, software configuration, implementation details, optimization parameters, etc.
5.1 Plant model and uncertain parameters
We consider a spacecraft, composed of one rigid body connected to two large flexible solar panels. Uncertainties are considered on various mechanical properties; in total, there are 43 uncertain parameters $\delta_{i}$. The spacecraft model $G(s,\delta)$ is generated with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox of [2, 35]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. Noting $M_{P}(\delta)$ the mass matrix at attachment point $P$ of a solar panel, and $L_{P}(\delta)$ the matrix of modal participation factors, the robustness analysis should exclude any configuration $\delta$ such that the residual mass matrix defined in Eq. (18) is not positive definite (indeed, such configurations do not represent a physical system). Hence the nonlinear constraint:
$M_{r}(\delta):=M_{P}(\delta)-L_{p}^{T}(\delta)L_{p}(\delta)\succeq 0$ (18)
which can also be written under a scalar form:
$c(\delta):=\alpha(-M_{r}(\delta))\leq 0$ (19)
and is assumed to be differentiable, at least close to worst-case configurations (which is often legitimate, as deficiency usually occurs along a single principal direction, leading to a simple maximum eigenvalue). The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (19).
5.2 Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite, and the output vector contains the angle and rate measurements polluted by some noise $n$. The closed loop is represented in Fig. 2, where $d(s)$ represents actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
Figure 2: Benchmark: closed-loop system
The control problem consists in placing a control bandwidth of 0.1 rad/s while ensuring a modulus margin of 0.5 and minimizing the variance of the actuator command in response to measurement noise. It reads:
[…]
where $\mathcal{D}$ is defined in (5) with the nonlinear constraint (19), the factor $\frac{1}{2}$ in the 2nd requirement ensures the desired modulus margin, and $W=\frac{10\mathrm{s}+1}{2(10\mathrm{s}+0.01)}I_{3}$ is a weighting filter to impose the desired bandwidth of 0.1 rad/s. The last requirement of problem (20) ensures closed-loop stability with some tolerance, where $A(K,\delta)$ is the matrix A of the state-space representation of the closed-loop system.
5.3 Worst-case search
In this section, we illustrate the worst-case search for the spectral abscissa, as presented in Section 3. We selected a controller $K_{1}$ obtained at an initial iteration of the robust controller tuning presented in Section 5.4, and analyze the robustness of the closed-loop system. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
SQP alone
We ran the SQP algorithm implemented in the Matlab routine fmincon [^1] As termination criteria, we set $10^{-6}$ as optimality tolerance (KKT stationarity condition), and $10^{-8}$ as step tolerance (step size between two iterations). from 100 random initial points. The resulting worst-case $\mathcal{H}_{\infty}$ norms of each run are in Fig. 3(a), and show the presence of several local optima, which highlights the limitation of such a local method. The runs took up to 51s (average: 4.7s).
Monte-Carlo + SQP
To show the benefit of initializing SQP after a Monte-Carlo search rather than from any random initial point, we repeated 100 times the method of Section 3 with $\epsilon=\gamma=0.01$: drawing $N=460$ samples, and selecting the worst-case point to initialize an SQP run. Fig. 3(b) presents the worst case for each group (in blue) and the worst case after the subsequent SQP (in red). Overall, MC+SQP performs consistently better than SQP alone or MC alone. Each group took about 0.57s (the samples were evaluated in parallel); the SQP runtimes were between 3s and 9s. Thus, the Monte-Carlo step comes at little to no extra cost (if parallel computing is available), and reduces the risk of falling into local minima.
(a) SQP
(b) MC and MC+SQP
Figure 3: Worst-case search results (spectral abscissa)
5.4 Robust control
In this section, we solve the robust control Problem (20) using Algorithm 2. We used the tolerances $\epsilon_{1}=5\%$ and $\epsilon_{2}=1\%$ (cf. Eq. (17)). The worst-case searches of Steps 2 and 3 were carried with $\epsilon=\gamma=0.01$ in Eq. (16) yielding $N=460$ Monte-Carlo samples, and Step 4 with $\epsilon=\gamma=0.001$ yielding $N=6910$ samples. We used $M=10$ (cf. the discussion following Eq. (16)).
For testing purpose, since the algorithm involves randomness from both Monte Carlo sampling, Algorithm 2 was executed 10 times. All runs were initialized with the same nominal controller, and all converged to a robust controller. Table 1 presents a few metrics. In particular, the most costly phase is the controller optimization (Step 1). Fig. 4 presents the transfer functions of Problem (20) after optimization of the controller. Active configurations (in red) include the worst cases detected in the final iteration.
Table 1: Controller optimization results
Metrics Min. Average Max.
Nb of iterations 6 8.4 12
Nb of active configurations 12 16.3 24
Total runtime (s) 1312 2257 3954
… including Step 1* runtime 66% 74% 85%
* Controller optimizaton (systune), as opposed to the worst-case searches (MC+SQP, Steps 2, 3 and 4)
Figure 4: Transfer functions in Problem (20) after controller optimization, with active configurations (red), 100 random samples (blue) and requirement (dashed lines)
6 Conclusion
This paper focused on the robust control and worst-case search in presence of nonlinear constraints restricting the hypercube. This specificity is missing from the current state of the art, and arises for example in mechanical systems.
Relying on the theory of upper-$C^{1}$ functions, we established favorable properties that justify the use of smooth optimization algorithms for these nonsmooth problems: any subgradient defines a valid descent direction, local optima satisfy the Karush–Kuhn–Tucker (KKT) conditions for all subgradients, and Sequential Quadratic Programming converges to KKT points when applied to such functions. This theoretical foundation legitimizes the use of standard, readily available optimization tools, which is of practical importance for industrial users.
To perform the worst-case search, we combined Monte-Carlo sampling for global exploration and Sequential Quadratic Programming for local exploitation. Numerical experiments on a flexible spacecraft benchmark show that this method is able to reliably detect even rare worst-case configurations. Then, by iteratively constructing a set of active configurations, multi-objectives robust controller optimization was performed in reasonable runtime.
Finally, the proposed worst-case search, performing global exploration (albeit without deterministic guarantee of finding the global optimum) and being compatible with a large number of uncertain parameters and states (as a gradient-based method), has the potential to complement existing robust control techniques (systune, $\mu$-analysis) even in the unconstrained case.
{ack}
The authors thank Vincent Guibert for his valuable feedback on the preprint.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT (GPT-5) to assist in improving the wording and reformulation of selected sentences. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Funding Sources
This work was funded par Région Occitanie in the frame of the France 2030 program. Ce projet a été financé par la Région Occitanie dans le cadre de France 2030.
Appendix A Mathematical proofs
In this Appendix, we prove the propositions of Section 2 that result from mere adaptations of existing literature. For readability, they are reminded before introducing the proofs.
Proposition 3 : Descent directions and optimality of upper-$C^{1}$ functions – Constrained case
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$. Let the $c_{i}$ be continuously differentiable.
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$ with $g_{x}\in\partial f(x)$ any subgradient. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$.
b) Optimality. (proved in the body of the paper).
{pf}
The proof of a) is an adaptation of [27, Theorem 3.1], originally written for a differentiable $f$, to the upper-$C^{1}$ case. Let us define:
$I=\{i:c_{i}(x)>0\}$
$\bar{I}=\{i:c_{i}(x)=0\}$
$\hat{I}=\{i:c_{i}(x)<0\}$
As in [27] we have:
$\theta^{\prime}_{r}(x,p)=f^{\prime}(x,p)+r\sum_{i\in I}\nabla c_{i}(x)^{T}p+r\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}\;.$
Since $f$ is upper-$C^{1}$, from [25, Proposition 2] we have $f^{\prime}(x,p)=\min_{g\in\partial f(x)}g^{T}p\leq g_{x}^{T}p$ and thus
$\theta^{\prime}_{r}(x,p)\leq g_{x}^{T}p+r\sum_{i\in I}\nabla c_{i}(x)^{T}p+r\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}\;.$
In the original proof of [27, Theorem 3.1], the expression above is an equality with $\nabla f(x)$ instead of $g_{x}$. From there, the proof of [27, Theorem 3.1] is identical, as follows. Since $(p,u$) is a KKT pair of $Q(x,H,g_{x})$, we have $c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0$ which yields $\nabla c_{i}(x)^{T}p\leq 0$ for $i\in\bar{I}$ and thus:
$\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}=0$
and since $u_{i}(c_{i}(x)+\nabla c_{i}(x)^{T}p)=0$ (again because $(p,u$) is a KKT pair of $Q(x,H,g_{x})$), we deduce:
$\theta^{\prime}_{r}(x,p)\leq g_{x}^{T}p+\sum_{i=1}^{n}u_{i}\nabla c_{i}(x)^{T}p\\ +\sum_{i=1}^{n}u_{i}c_{i}(x)+r\sum_{i\in I}\nabla c_{i}(x)^{T}p\;.$
Since $(p,u$) is a KKT pair of $Q(x,H,g_{x})$ we also have:
$g_{x}+\sum_{i=1}^{n}u_{i}\nabla c_{i}(x)+\frac{1}{2}(H+H^{T})p=0$
and noticing that $\sum_{i\in\bar{I}\cup\hat{I}}^{n}u_{i}c_{i}(x)\leq 0$ yields:
$\theta^{\prime}_{r}(x,p)\leq-\frac{1}{2}p^{T}(H+H^{T})p+\sum_{i\in I}(u_{i}c_{i}(x)+r\nabla c_{i}(x)^{T}p)$
and then, using again that $c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0$:
$\theta^{\prime}_{r}(x,p)$ $\leq-\frac{1}{2}p^{T}(H+H^{T})p+\sum_{i\in I}(u_{i}-r)c_{i}(x)$
$<0$
since $H$ is positive definite and $|u|_{\infty}<r$.
Lemma 2: Perturbation of a quadratic problem
Let us consider the two quadratic problems:
$x_{0}=\arg\min\{\;Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\;\;\\ \text{subject to }\;Gx\leq g\}$
$x_{0}^{\prime}=\arg\min\{\;Q^{\prime}(x)=\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\;\;\\ \text{subject to }\;G^{\prime}x\leq g^{\prime}\}$
[…]
$x_{0}^{\prime}-x_{0}||\leq c\epsilon$
whenever $\epsilon<\epsilon_{0}$.
{pf}
[…]
$x_{0}=\arg\min\left\{Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\right\}$
$x^{\prime}_{0}=\arg\min\left\{Q^{\prime}(x)\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\right\}$
we have:
[…]
The proof for this statement is almost identical to the one provided in [22, Theorem 2.1], as follows. For any $x\in\mathbb{R}^{n}$, we have:
$(x-x_{0})^{T}\nabla Q(x_{0})\geq 0$
$(x-x^{\prime}_{0})^{T}\nabla Q^{\prime}(x^{\prime}_{0})\geq 0$
with the gradients $\nabla Q(x)=Kx-k$ and $\nabla Q(x)=K^{\prime}x-k^{\prime}$. Substituting $x=x^{\prime}_{0}$ and $x=x_{0}$ respectively in these two equations and adding the results yields
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q(x_{0})-\nabla Q^{\prime}(x^{\prime}_{0}))\geq 0$
which is equivalent to
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q^{\prime}(x^{\prime}_{0})-\nabla Q^{\prime}(x_{0}))\\ \leq(x^{\prime}_{0}-x_{0})^{T}(\nabla Q(x_{0})-\nabla Q^{\prime}(x_{0}))$
For the left side:
[…]
For the right side:
[…]
Then, the rest of the proof provided in [22] still holds. Indeed, the propositions [22, Propositions 3.1, 3.2] only tackle the constrained sets, and do not involve $k$ or $k^{\prime}$. Moreover, in, [22, Section 4], the relationship between $k$ and $k^{\prime}$ is not used (except when using [22, Theorem 2.1], which we generalized already); in fact, [22, Section 4] only works on $Q(x)$.
Proposition 7 : Convergence of SQP for upper-$C^{1}$ functions
Let $f$ be upper-$C^{1}$ and the $c_{i}$ be continuously differentiable. Assume that the following conditions are satisfied:
(i) there exists $\alpha,\beta>0$ such that, for each $k$ and for any $x\in\mathbb{R}^{n}$:
$\alpha x^{T}x\leq x^{T}H_{k}x\leq\beta x^{T}x$
(ii) For each $k$, there exists a KKT point of $Q(x_{k},H_{k},g_{x_{k}})$ (equation (10)) with a Lagrange multiplier vector $u_{k}$ such that $u_{k}||_{\infty}\leq r$.
Then, any sequence $(x_{k})$ generated by Algorithm 1 either terminates at a KKT point of the constrained optimization problem (8), or any accumulation point $\bar{x}$ with
$S^{0}(\bar{x})=\{p:c(\bar{x})+\nabla c(\bar{x})^{T}p<0\}\neq\emptyset$
is a KKT point of Problem (8).
{pf}
Once again, we follow the proof of [27, Theorem 3.2] and adapt it to the case where f is not differentiable but only upper-$C^{1}$. For completeness, we reproduce the entire proof and indicate where our modified assumptions apply, while emphasizing that the proof itself is not our own.
By assumption (ii), $(p_{k},u_{k})$ is a KKT pair of $Q(x_{k},H_{k},g_{x_{k}})$ with $u_{k}||_{\infty}\leq r$. If $p_{k}=0$, then $(x_{k},u_{k})$ satisfies the KKT conditions (9) of Problem (8) and the algorithm can terminate at $x_{k}$.
Thus, for the following, we assume that $p_{k}\neq 0$. From Proposition 3.a) (originally [27, Theorem 3.1], which we adapted to upper-$C^{1}$ functions), it is possible to choose $x_{k+1}$ as described in Algorithm 1, and in particular, it verifies:
$\theta_{r}(x_{k+1})<\theta_{r}(x_{k})+\epsilon_{k}\;.$
Let $\bar{x}$ be an accumulation point of $(x_{k})$ with $S^{0}(\bar{x})\neq\emptyset$. Without loss of generality (passing to the subsequence if necessary) we may assume
$x_{k}\rightarrow\bar{x}\;,H_{k}\rightarrow\bar{H}$
where the existence of $\bar{H}$ follows from assumption (i).
In the original proof [27], the continuity of the gradient of $f$ ensures that $\nabla f(x_{k})$ converges to $\nabla f(\bar{x})$. Here, a similar property is maintained as follows. Since $f$ is locally Lipschitz, its subgradients $(g_{x_{k}})$ are uniformly bounded for $k$ sufficiently large, hence there exists a subsequence that converges to a vector $\bar{g}$. Without loss of generality (passing to the subsequence if necessary) we note
$g_{x_{k}}\rightarrow\bar{g}\;,$
and the upper semi-continuity of the Clarke subdifferential [19, Proposition 1.5] yields $\bar{g}\in\partial f(\bar{x})$. Since $\bar{H}$ is positive definite and $S^{0}(\bar{x})\neq\emptyset$, the quadratic problem $Q(\bar{x},\bar{H},\bar{g})$ (originally $Q(\bar{x},\bar{H},\nabla f(\bar{x}))$ in [27]) has a unique KKT point $\bar{p}$. If $\bar{p}=0$, then $\bar{x}$ is a KKT point of Problem (8) and the theorem follows. Thus, we now assume that $\bar{p}\neq 0$ and we will show that this leads to a contradiction.
From the continuity of $\nabla c(x)$, and by Lemmas 1 and 2, where $x_{k+1}$ and $x_{k}$ play the roles of $x_{0}$ and $x^{\prime}_{0}$, and $g_{x_{0}}$ and $g_{x^{\prime}_{0}}$ play the roles of $k$ and $k^{\prime}$, we deduce that:
$p_{k}\rightarrow\bar{p}$
In the original proof [27], this follows [22] (the equivalent to our Lemma 2) and the continuity of $\nabla f(x)$ instead (which is stronger than our Lemma 1).
Since $(u_{k})$ is uniformly bounded by $r$, it has an accumulation point $\bar{u}$. We note
$u_{k}\rightarrow\bar{u}$
with $\bar{u}||_{\infty}\leq r$. By passing to the limit in the KKT relation, which is valid for any $k$, we obtain that $\bar{u}$ is a Lagrange multiplier of $Q(\bar{x},\bar{H},\bar{g})$.
The rest of the proof of [27] is unchanged, and is as follows. Let $\bar{\lambda}\in\begin{bmatrix}0&\lambda_{max}\end{bmatrix}$ be chosen such that
$\theta_{r}(\bar{x}+\bar{\lambda}\bar{p})=\min_{0\leq\lambda\leq\lambda_{max}}\;\theta_{r}(\bar{x}+\lambda\bar{p}).$
Let us set
$\gamma=\theta_{r}(\bar{x}+\bar{\lambda}\bar{p})-\theta_{r}(\bar{x})>0$
which is strictly positive by Proposition 3.a). Since $x_{k}+\bar{\lambda}p_{k}\rightarrow\bar{x}+\bar{\lambda}\bar{p}$, for sufficiently large $k$ we have:
$\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2<\theta_{r}(\bar{x})\;.$
However, for all $k$ we have $\theta_{r}(x_{k+1})<\theta_{r}(x_{k})+\epsilon_{k}$, and for sufficiently large $k$ we have
$\sum_{i=k}^{\infty}\epsilon_{i}<\gamma/2$
which leads to:
$\theta_{r}(\bar{x})$ $<\theta_{r}(x_{k+1}+\sum_{i=k+1}^{\infty}\epsilon_{i}$
$\leq\min_{0\leq\lambda\leq\lambda_{max}}\theta_{r}(x_{k}+\lambda p_{k})+\epsilon_{k}+\sum_{i=k+1}^{\infty}\epsilon_{i}$
$<\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2$
which contradicts the previous inequality $\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2<\theta_{r}(\bar{x})$. This proves that $\bar{p}=0$ and that $\bar{x}$ is a KKT point of Problem (8).